OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 6.

OpenStudy (anonymous):



OpenStudy (anonymous):

A. y = -2x2 + 2x -2; –58 B. y = 2x2 - 2x -2; 60 C. y = 2x2 - 2x -2; 58 D. y = -2x2 + 2x +2; –58

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

it takes a minute, but you could just try those x and y values in those 4 equations to find the one that works.. There's another way, but this way might be more appropriate under the circumstances :)

OpenStudy (anonymous):

A. y = -2x2 + 2x -2; –58 B. y = 2x2 - 2x -2; 60 C. y = 2x2 - 2x -2; 58 D. y = -2x2 + 2x +2; –58

OpenStudy (anonymous):

um how do i put these in my calculator in the table

OpenStudy (anonymous):

I don't know... don't have your calculator. Just "say" these out loud with me... :) (don't have to say them... just do them in your head... you can do it...) Take the first point, x = 1, y = 2... try the equation in answer A: A. y = -2x2 + 2x -2; –58 so -2(1)^2 + 2(1) - 2 = -2 + 2 - 2 = -2... but it's supposed to be y = 2, so this equation doesn't work. Try B next...

OpenStudy (anonymous):

ummm uhhhh lol i got lost on A with all that work

OpenStudy (anonymous):

Or, an even better first start is to realize that for x = 0, y = -2. So ignore all the x-term parts of the equation except the +2 or -2 at the ends, since for x = 0, those disappear. (all the x terms = 0, since x = 0). So the part at the end has to be a -2... that rules out D, at least...

OpenStudy (anonymous):

B: 2(1)^2 - 2(1) - 2

OpenStudy (anonymous):

b isnt 60

OpenStudy (anonymous):

yep nope im confused

OpenStudy (anonymous):

No, those numbers like 58 and 60 would be your predictions using this formula for when x = 6, for the 2nd part of the problem... confusing is right though :)

OpenStudy (anonymous):

im about to guess cause i just screwed up my whole thinking

OpenStudy (anonymous):

Is there any chance you mixed up that table of numbers? I haven't found any of A - D that actually works...

OpenStudy (anonymous):

i wish i messed it up lol, what i copied over is what the test says A. y = -2x^2 + 2x -2; –58 B. y = 2x^2 - 2x -2; 60 C. y = 2x^2 - 2x -2; 58 D. y = -2x^2 + 2x +2; –58

OpenStudy (anonymous):

I looked at it wrong... sorry :(

OpenStudy (anonymous):

I suspect help is here though :)

OpenStudy (anonymous):

lol yea i started this test and was like uuuummmmmmm yea right im supposed to pass this hahah lol

OpenStudy (anonymous):